mikewwe13:

What is the average rate of change of the function f(x)=4(2)^x from x = 0 to x = 2? Enter your answer in the box. ___________________

sillybilly123:

compute as follows: \(\dfrac{\text{change in f(x)}}{\text{change in x}} = \dfrac{ f(x_2) - f(x_1)}{x_2 - x_1} = \dfrac{4(2)^2 - 4(2)^0 }{2- 0 }\)